A dog (mass 11.7 kg) sits on a sled (mass 5.1 kg) that is being pushed with a force of 35 N on a flat snowy surface. If the acceleration of both is 1.7 m/s2, what is the coefficient of kinetic friction for the sled sliding on the snow?
1
Expert's answer
2019-03-21T10:54:08-0400

The Newton's second law states

"ma=F-F_{\rm{fric}}"

"(11.7+5.1)1.7=35-F_{\rm{fric}}"

So, the friction force

"F_{\rm{fric}}=6.44\:\rm{N}"

The coefficient of kinetic friction

"\mu_k=\frac{F_{\rm{fric}}}{mg}=\frac{6.44}{(11.7+5.1)9.81}=0.04"